OpenStudy (anonymous):

Multiply and simplify. y^2-6y-27/y^2-6y+9 * y-3/y-9 Can someone please help me solve this, and show me how to do the work, so I can do the rest myself? Thank you.

OpenStudy (anonymous):

factor each part and then cancel equivalent sections y^2 -6y-27 = (y-9)(y+3) y^2 -6y +9 = (y-3)(y-3) Now just cancel :D

OpenStudy (anonymous):

Thank you. :)